how much pure acid should be mixed with 2 gallons of a 40% acid solution in order to get a 70% acid solution
Answer by Earlsdon(6294)

Let x = the number of galons of pure acid to be added to the 2 gallons of 40% acid solution to obtain (2+x) gallons of 70% acid solution. You can express this algebraically as: (40% = 0.4) and (70% = 0.7)
x + 2(0.4) = (2+x)(0.7) Simplify and solve for x.
x+0.8 = 1.4+0.7x Subtract 0.7x from both sides.
0.3x+0.8 = 1.4 Now subtract 0.8 from both sides.
0.3x = 0.6 Finally, divide both sides by 0.3
x = 2
You will need to mix 2 gallons of pure acid to the 2 gallons of 40% acid solution to obtain 4 gallons of 70% acid solution.